the potential energy of an objectย  of mass m moving in xy plane in a conservative field is given by U=ax+by, where x and y are positionย  coordinates of theย  object. find magnitude of its acceleration ?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Fx = โ€“ dU/dx = > -(a)
Fy = โ€“ dU/dy => -(b)
maccln.x = -(a)ย 
or
accln.x = -a/m
and
maccln.y = -(b)
or
accln.y = -b/m
Net accln. = root((a/m)^2 + (b/m)^2)